You have a box that is a good size for your tape collection. Two rolls of tape’s will fit in the box. The box is 10 inches wide. Each tape is 5/8 inches wide. How many tables will fit in the box?
1
Expert's answer
2019-09-02T12:44:44-0400

Each roll contains no more than"\frac{10}{\frac{5}{8}}=\frac{10\cdot 8}{5}=16" tapes. Since 16 is integer, we have that each roll contains 16 tapes.

So 2 rolls contain "16\cdot 2=32" tapes.